The perimeter of a rectangle is 78cm. If the length is six more than twice the width, find the dimensions of the rectangle.
Answer by AHBK(13)

let W=width and L=length of rectangle
perimeter=78
2W+2L=78
and we have the length is six more twice the width so:
L=2w+6
subtitute 2W+6 instead of L above
so we have
2W+2(2W+6)=78
2W+4W+12=78
6W=78-12
6W=66
W=66/6
W=11cm
now substitute W=11cm in 2W+2L=78
so
2(11)+2L=78
22+2L=78
2L=78-22
2L=56
L=56/2
L=28cm
your welcome